# LightGaussian: Unbounded 3D Gaussian Compression with 15x Reduction and 200+ FPS

### v5

@@ Abstract
Recent advancements in real-time neural rendering using point-based techniques have paved the way for the widespread adoption of 3D representations. However, foundational approaches like 3D Gaussian Splatting come with a substantial storage overhead caused by growing the SfM points to millions, often demanding gigabyte-level disk space for a single unbounded scene, posing significant scalability challenges and hindering the splatting efficiency. To address this challenge, we introduce LightGaussian, a novel method designed to transform 3D Gaussians into a more efficient and compact format. Drawing inspiration from the concept of Network Pruning, LightGaussian identifies Gaussians that are insignificant in contributing to the scene reconstruction and adopts a pruning and recovery process, effectively reducing redundancy in Gaussian counts while preserving visual effects. Additionally, LightGaussian employs knowledge distillation and pseudo-view augmentation to transfer spherical harmonics efficients to a lower degree, allowing knowledge convert to more compact representations while maintaining scene appearance. Furthermore, we propose a Gaussian Vector Quantization based on Gaussian global significance, to quantize all redundant attributes, resulting in lower bitwidth representations with minimal accuracy losses. In summary, LightGaussian achieves an averaged compression rate over 15$\times$ while boosting the FPS from 119 to 209, enabling an efficient representation of complex scenes on Mip-NeRF 360, Tank & Temple datasets.
Figure 1: Compressibility and Rendering Speed. We present LightGaussian to transform 3D Gaussians into a more compact representation. LightGaussian effectively prunes redundant Gaussians while preserving visual fidelity (on the left). It reduces the average storage from 724MB to 42MB and improves the FPS from 119 to 209.
@@ 1 Introduction
Novel view synthesis (NVS) aims to generate photo-realistic images of a 3D scene from unobserved viewpoints, given a set of calibrated multi-view images. NVS holds importance as it can be used for a wide range of real-world applications, including virtual reality [12], augmented reality [71], digital twin [13], and autonomous driving [64]. Neural Radiance Fields (NeRFs)[45, 4, 5] have demonstrated promising ability for photo-realistic 3D modeling and synthesis from multi-view images where 3D location and view directions are mapped to view-dependent color and volumetric density. The pixel intensity can be rendered using the volume rendering technique[14]. However, NeRF and its variants have struggled with rendering speeds for practical deployment in real-world scenarios. Subsequent efforts along this trend introduce voxel-based implicit representation [57, 62, 54, 67, 10], hash grid [46], computational parallelism [51], or neural light field representation [61] to speed up the rendering. However, they either require specific design or face a trade-off between rendering quality and efficiency, making them difficult to generalize to large-scale scenarios with practical speed. Recent progress in point-based 3D Gaussian Splatting (3D-GS) [32] has brought photo-realistic rendering quality to the real-time level, even for complex scenes. The authors utilize the idea of representing the scene as explicit 3D Gaussians with specific properties to model the scene; the 2D images are efficiently rendered using a technique named splatting[36]. The optimal balance between speed and quality indicates a potential trend to use 3D-GS as a new representation for generating numerous large-scale scenes in digital twins and autonomous driving. However, using point-based representations inevitably introduces significant storage costs, as each point and its attributes are stored independently. For instance, 3D-GS on a typical unbounded 360 scene [5] requires more than one gigabyte, which inhibits the scalability of 3D-GS (e.g., 1.4GB on scene Bicycle).
In this paper, we counteract the heavy storage issue and deliver a compact representation while preserving the rendering quality. By examining the well-trained point-based representation, each scene is composed of millions of Gaussians, grown from sparse Structure from Motion (SfM) point cloud. Attributes are attached to Gaussians to model the scene’s geometry and appearance. However, the significant number of Gaussians after its optimization, as well as the high-degree Spherical Harmonics (SH) coefficients used for modeling scene reflection, contribute to an over-parametrized representation when fitting to the scenes. To minimize the required Gaussian number, we propose proper criteria to measure the global significance of each 3D Gaussian, in the context of its contribution for view synthesis. Gaussians that have minimal impact on visual quality are identified and pruned. This is followed by streamlined recovery steps that allow for rapid optimization, named as Gaussian Pruning and Recovery. Spherical harmonics (SH) coefficients, constituting a significant portion of the data, are employed to model view-dependent color and to capture the interaction of light with surfaces from diverse angles. However, utilizing higher-degree SH coefficients presents a trade-off between accuracy and computational efficiency. While higher-degree coefficients offer more detailed lighting information, enhancing scene realism, they also escalate computational complexity and storage demands. To mitigate this, we propose a SH Distillation step complemented by pseudo-view augmentation. This approach is designed to efficiently condense the information from higher-degree coefficients into a more compact form, ensuring that quality is preserved without inflating computational requirements. Besides focusing on reducing redundancy in spatial (Gaussian count) and lighting (SHs) representations, we additionally introduce a Vector Quantization step. This process adaptively selects a specific quantity of distinct point attributes, determined by their global significance, which further reduces the required bitwidth from the original format.
In summary, our proposed framework, LightGaussian, efficiently reduces the Gaussian count (e.g., from 1.49M to 575K, Fig. 1, left), significantly reducing storage requirements from 724MB to 42MB, with minimal rendering quality decrease ($\downarrow$ 0.005 in SSIM) on the Mip-NeRF 360 dataset (Fig. 1, right). LightGaussian further improves the rendering speed to a higher level (200+ FPS) on complex scenes containing detailed backgrounds, suggesting a viable solution for broadening the application scope.
@@ 2 Related Works
Efficient 3D Scene Representations for NVS.
Neural radiance field
(NeRF) [45] propose to use multi-layer perceptron (MLPs) to represent a scene, and this compact representation has brought view synthesis quality to a new stage. However, NeRF suffers from extremely slow inference challenges. Follow-ups either explore ray re-parameterizations[4, 5], explicit spatial data structures [68, 21, 41, 31, 57, 46, 10], caching and distillation [61, 22, 28, 51], or ray-based representations [2, 56] for speeding up. Still, NeRF-based methods struggle to achieve real-time rendering speed in practical large-scale scenes, which is caused by the multiple queries needed for rendering a single pixel, limiting their practical use. Point-based representations (e.g., Gaussian) have been extensively explored in various domains, including shape reconstruction [33], molecular structures modeling [7], and the replacement of point-cloud data [15]. Additionally, these representations are utilized in shadow [49] and cloud rendering applications [43]. Recently, Pulsar [37] has shown its efficient sphere rasterization, 3D Gaussian Splatting (3D-GS) [32] utilizes anisotropic Gaussians [72] and introduces a tile-based sorting for rendering, achieves real-time speed with comparable rendering quality to the MLP-based ones(e.g., Mip-NeRF 360 [5]). Although promising, the heavy storage requirement to store all attributes attached to the Gaussians often requires gigabyte-level disk space for saving a single unbounded scene, and millions of Gaussians hinder the rendering efficiency of 3D-GS. Concurrent works adopt region-based vector quantization [48], K-means based codebooks [47], or learned binary masks for each Gaussian, replacing the SHs with grid-based neural networks [39] within a modified training framework to reduce model size.
Model Pruning and Vector Quantization.
Model pruning involves reducing the complexity of a neural network by eliminating non-significant parameters to maintain a balance between performance and resource utilization. Unstructured [38] and structured pruning [27, 1] remove components at weight-level and neuron (mostly channel) levels to provide a smaller network with a smaller or more efficient network architecture. The iterative magnitude pruning (IMP) method, where weights of the smallest magnitude are progressively pruned over multiple iterations, has been highly successful in lottery ticket rewinding [19, 20]. Additionally, vector quantization [59] aims to represent data with discrete entries of a learned codebook (i.e., tokens) to achieve lossy compression. In general, the mean square error (MSE) is used to find the most similar pattern in the codebook to replace the original input data vector. Previous works [11, 23, 44] have shown that learning a discrete and compact representation not only contributes to visual understanding but also improves the robustness of models. In this vein, vector quantization has been widely adopted in image synthesis [16], text-to-image generation [24], and novel view synthesis [68, 40, 25].
@@
Knowledge Distillation (KD) is widely adopted in various machine learning tasks [55, 53, 63, 61]. A student model is trained for the primary purpose of model compression [8] or ensembling [30] with the help of teacher models. Many variants [3, 52, 60, 69, 29, 42, 58] have been proposed to achieve better knowledge transfer from the teacher to the student models. In the field of 3D vision, neural scene representations have embraced knowledge distillation mainly through view renderings to leverage existing 2D priors. DreamFusion [50] and NeuralLift-360 [65] adopt pre-trained text-to-image diffusion models for 3D generation. DFF [35], NeRF-SOS [18], INS [17], SA3D [9] distill various pre-trained 2D image feature extractors to perform corresponding tasks in the 3D domain. Knowledge distillation has also played a key role in model compression of scene representations. R2L [61] and KiloNeRF [51] distill Neural Radiance Fields (NeRF) into more efficient representations such as light fields or multiple tiny MLPs. Our work falls into the category of utilizing knowledge distillation for model compression and leverages novel view renderings as the bridge to connect the teacher and student models.
@@ 3 Methods
Figure 2: The overall pipeline of LightGaussian. 3D Gaussians are optimized from multi-view images and SfM points. LightGaussian first calculates the global significance for each Gaussian on training data, where Gaussians with the least significance are pruned. A subsequent distillation with synthesized pseudo-views is introduced to transfer the SH into a compact format. Vector quantization, consisting of codebook initialization and assignment, is further introduced for reducing the model bandwidth.
@@
The overview of LightGaussian is illustrated in Fig. 2. The 3D-GS model is trained using multi-view images and is initialized from SfM point clouds. By expanding the sparse points to millions of Gaussians, the scene is well-represented. Then, the 3D-GS undergoes processing within our pipeline to transform it into a more compact format. This involves utilizing Gaussian Prune and Recovery to reduce the number of Gaussians, SH Distillation to remove redundant SHs while preserving the essential lighting information, and Vector Quantization to store Gaussians at a lower bit-width.
@@ 3.1 Background: 3D Gaussian Splatting
where $\boldsymbol{\Sigma}$ can be decomposed into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ for differentiable optimization.
@@ 3.2 Gaussian Pruning & Recovery
Gaussian densification [32], which involves cloning and splitting the initial SfM point cloud, is employed to address the challenge of insufficient coverage and is used to model small-scale geometry as well as detailed scene appearance. While this strategy leads to significantly improved reconstruction quality, it results in the number of Gaussians growing from thousands to millions after optimization. Such an explicit point-based representation with a large number of Gaussians requires an extremely significant storage overhead. Drawing inspiration from the success of representative neural network pruning techniques [26], which eliminates less impactful neurons without compromising the network’s overall performance, we tailor a general pruning paradigm for 3D Gaussian Splatting representation to reduce the over-parameterized point number in a manner that preserves the original accuracy. Therefore, identifying the most representative redundant Gaussians with recoverable accuracy is a crucial step in our approach. However, pruning the Gaussians based on simplistic criteria (e.g., point opacity) can lead to a substantial degradation in modeling performance, especially where the intricate scene structure may be eliminated, as is demonstrated in Fig. 3.
@@
Rather than merely depending on Gaussian opacity for assessing significance, we are inspired by Eq. 2. In this context, 3D Gaussians within the view frustum are evaluated and rasterized by projecting them onto a camera viewpoint for rendering. The initial significance score of each Gaussian (denoted as $\text{G}(\boldsymbol{X}_{j})$) can then be quantified based on its contribution to each pixel (ray, $r_{i}$) across all training views using the criteria $\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})$ whether they intersect or not. Consequently, we iterate over all training pixels to calculate the hit count of each Gaussian. The score is further refined by the adjusted 3D Gaussian’s volume $\large{\gamma}$$(\boldsymbol{\Sigma}_{j})$ and 3D Gaussian’s opacity $\sigma_{j}$, as they all contribute for the rendering formula. The volume calculation equation of the $j$-th Guassian is V${(\boldsymbol{\Sigma}_{j})}=\frac{4}{3}\pi abc$, where $abc$ are the 3 dimensions of Scale ($\mathbf{S}$). Finally, the global significant score can be summarized as:
$\text{GS}_{\text{j}}$ $=\sum_{i=1}^{MHW}\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})\cdot\sigma_{j}\cdot{\large{\gamma}}(\boldsymbol{\Sigma}_{j}),$ (3)
where $j$ is the Gaussian index, $i$ means a pixel, $M$, $H$, and $W$ represents the number of training views, image height, and width, respectively. $\mathbbm{1}$ is the indicator function that determines whether a Gaussian intersects with a given ray.
However, the use of Gaussian volume tends to exaggerate the importance of background Gaussians, leading to the immoderate pruning of Gaussians that model intricate geometry. Therefore, we propose a more adaptive way to measure the dimension of its volume:
@@
$\text{V}_{\text{norm}}$ $=\text{min}\left(\text{max}\left(\frac{\text{V}(\boldsymbol{\Sigma})}{\text{V}_{\text{max90}}},0\right),1\right).$
@@
We rank all the Gaussians based on the computed global significance scores, which provides a quantitative basis for pruning Gaussians with lower scores. The remaining ones are optimized (adapted) to jointly adjust their attributes, without further densification, to compensate for the slight loss incurred by pruning. Note that, we utilize photometric loss on original training views with 5,000 iterations which is aligned with 3D-GS training.
@@ 3.3 Distilling into Compact SHs
In the uncompressed Gaussian Splat data, a considerable portion is occupied by Spherical Harmonics (SH) coefficients, necessitating (45+3) floating-point values per splat, which account for 81.3 percent of the total attribute volume. While directly diminishing the degree of SH is advantageous for reducing disk space consumption, this action results in a significant reduction in surface “shininess”, notably impacting the variance in specular reflections as the viewpoint changes.
To strike a balance between model size and scene modeling quality, we suggest transferring knowledge from well-trained high-degree SHs to their more compact, lower-degree counterparts through a distillation process. Specifically, we distill the knowledge from the teacher model, which employs full-degree SHs and remains static, to the student model, characterized by truncated lower-degree SHs. The supervision is based on the differences in the predicted pixel intensities, denoted as $\boldsymbol{C}$, between the two models:
$\mathcal{L}_{\text{distill}}=\frac{1}{HW}\sum_{i=1}^{HW}\left\lVert\boldsymbol{C}_{\text{teacher}}(r_{i})-\boldsymbol{C}_{\text{student}}(r_{i})\right\rVert_{2}^{2}.$ (5)
Simply reintroducing these view-dependent visual effects to the network does not significantly enhance its knowledge. We, therefore, suggest employing the concept of data augmentation. This involves training the Gaussians not only to represent known views but also to learn behaviors from unseen (pseudo) views modeled by the teacher model, thus broadening their representational capacity.
Synthesize Pseudo Views.
Specular reflection rays are reflected off the surface of an object when the viewpoints are altered. To enrich the training data, we can introduce additional viewpoints through sampling, which are capable of capturing such reflectance characteristics. Given our knowledge of the camera locations and viewing directions for all training views, we strategically sample pseudo views around each training view, adhering to a Gaussian distribution, while maintaining consistent camera viewing directions:
@@
where $\mathbf{t}_{\text{pseudo}}$ and $\mathbf{t}_{\text{train}}$ represent the newly synthesized and training camera positions, respectively. $\mathcal{N}$ denotes a Gaussian distribution with mean 0 and variance $\sigma^{2}$, which is added to the original position to generate the new position.

### v6

@@ Abstract
Recent advances in real-time neural rendering using point-based techniques have enabled broader adoption of 3D representations. However, foundational approaches like 3D Gaussian Splatting impose substantial storage overhead, as Structure-from-Motion (SfM) points can grow to millions, often requiring gigabyte-level disk space for a single unbounded scene. This growth presents scalability challenges and hinders splatting efficiency. To address this, we introduce LightGaussian, a method for transforming 3D Gaussians into a more compact format. Inspired by Network Pruning, LightGaussian identifies Gaussians with minimal global significance on scene reconstruction, and applies a pruning and recovery process to reduce redundancy while preserving visual quality. Knowledge distillation and pseudo-view augmentation then transfer spherical harmonic coefficients to a lower degree, yielding compact representations. Gaussian Vector Quantization, based on each Gaussian’s global significance, further lowers bitwidth with minimal accuracy loss. LightGaussian achieves an average 15$\times$ compression rate while boosting FPS from 144 to 237 within the 3D-GS framework, enabling efficient complex scene representation on the Mip-NeRF 360 and Tank & Temple datasets. The proposed Gaussian pruning approach is also adaptable to other 3D representations (e.g., Scaffold-GS), demonstrating strong generalization capabilities.
   Project Website: https://lightgaussian.github.io
Figure 1: Compressibility and Speed: LightGaussian compacts 3D Gaussians, reducing storage from 782MB to 45MB and boosting FPS from 144 to 237, while maintaining visual quality.
@@ 1 Introduction
Novel view synthesis (NVS) aims to generate photo-realistic images of a 3D scene from unobserved viewpoints, given a set of calibrated multi-view images. This capability has widespread applications in virtual reality [1], augmented reality [2], digital twins [3], and autonomous driving [4]. Neural Radiance Fields (NeRFs) [5, 6, 7] have shown promise in 3D modeling from multi-view images by mapping 3D locations and view directions to view-dependent color and volumetric density, with pixel intensity rendered through volume rendering [8]. However, NeRF and its variants face limitations in rendering speed, limiting their deployment in real-world scenarios. To address this, voxel-based representations [9, 10, 11, 12, 13], hash grids [14], and neural light fields [15] have been developed. Despite improvements, these methods often compromise between quality and speed.
Recent progress in point-based 3D Gaussian Splatting (3D-GS) [16] has enabled real-time rendering with photo-realistic quality for complex scenes. By representing the scene with explicit 3D Gaussians and using a splatting technique [17], 3D-GS balances speed and quality, making it suitable for large-scale scenarios like digital twins and autonomous driving. However, point-based methods incur high storage costs, as each point and its attributes require independent storage. Additionally, heuristic densification of sparse SfM points into dense Gaussians often results in overparameterization, leading to excessive storage and slower rendering speeds. For instance, a typical unbounded 360-degree scene in 3D-GS [7] may require over 1GB of storage (e.g., 1.4GB for the Bicycle scene).
In this paper, we address storage and rendering speed issues by developing a compact representation that retains the original rendering quality. The heuristic densification process in 3D-GS results in significant redundancy. Our method, LightGaussian, reduces redundancy by targeting both Gaussian count (N) and feature dimension (F) through a comprehensive pipeline:
  • To reduce Gaussian count (N), we propose a Gaussian Pruning and Recovery step, identifying and removing Gaussians with minimal impact on visual quality, followed by recovery to ensure smooth adaptation.
  • For compressing features (F), we introduce an SH Distillation process to compact higher-degree spherical harmonic (SH) coefficients, supported by pseudo-view augmentation. Additionally, we employ Vector Quantization (VQ) to adaptively select a codebook of Gaussian attributes (e.g, positions, scales, and rotations), reducing precision for less significant features and applying quantization-aware fine-tuning to maintain quality.
In summary, LightGaussian achieves substantial compression (e.g., from 782MB to 45MB) while maintaining visual fidelity (SSIM decrease of only 0.007 on Mip-NeRF 360). Rendering speed also improves, reaching over 200 FPS on complex scenes. Our Gaussian Pruning and Recovery method generalizes well, enhancing performance across different 3D Gaussian formats, such as Scaffold-GS.
@@ 2 Related Works
Efficient 3D Scene Representations.
Neural radiance fields (NeRF) [5] use multi-layer perceptrons (MLPs) to represent scenes, setting new standards for view synthesis quality. However, NeRF models face challenges with slow inference speeds, limiting practical use. Efforts to address this have explored ray re-parameterizations [6, 7], explicit spatial data structures [18, 19, 20, 21, 9, 14, 13], caching and distillation techniques [15, 22, 23, 24], and ray-based representations [25, 26]. Nevertheless, achieving real-time rendering remains challenging for NeRF methods, especially for large-scale scenes where multiple queries per pixel hinder performance.
Point-based representations, like Gaussians, have been explored in various applications, such as shape reconstruction [27], molecular modeling [28], and point-cloud replacement [29], as well as in shadow [30] and cloud rendering [31]. Pulsar [32] demonstrated efficient sphere rasterization, while 3D Gaussian Splatting (3D-GS) [16] applies anisotropic Gaussians [33] with tile-based sorting to achieve real-time speed and quality comparable to MLP-based methods like Mip-NeRF 360 [7].
Despite its strengths, 3D-GS has high storage demands due to the extensive attributes stored with each Gaussian, often requiring gigabytes per scene and hindering rendering efficiency. Recent concurrent works address these issues by using region-based vector quantization [34], K-means codebooks [35], view-direction exclusion [36], or learned binary masks and grid-based neural networks instead of spherical harmonics (SHs) [37] to reduce model size.
Pruning and Quantization.
Model pruning reduces neural network complexity by removing non-significant parameters, balancing performance and resource use. Unstructured [38] and structured pruning [39, 40] eliminate parameters at the weight level and neuron/channel levels, resulting in a smaller, more efficient architecture. Iterative magnitude pruning (IMP), where low-magnitude weights are progressively removed, has proven effective in methods like lottery ticket rewinding [41, 42].
Vector quantization (VQ) [43] compresses data by representing it with discrete codebook entries (tokens), using mean square error (MSE) to select the closest match in the codebook for each data vector. Prior studies [44, 45, 46] have shown that learning discrete, compact representations enhances visual understanding and model robustness. VQ has thus been widely applied in image synthesis [47], text-to-image generation [48], and novel view synthesis [18, 49, 50].
@@
Knowledge Distillation (KD) [51, 52, 53, 54, 15] trains a smaller student model by transferring knowledge from a larger teacher model [55]. In 3D vision, KD has been applied to neural scene representations, leveraging view renderings to incorporate 2D priors. For example, DreamFusion [56] and NeuralLift-360 [57] use pre-trained diffusion models for 3D generation, while models like DFF [58], NeRF-SOS [59], INS [60], and SA3D [61] distill 2D image feature extractors to 3D tasks. KD has also been central to compressing scene representation models [15, 24].
@@ 3 Methods
Figure 2: Pipeline of LightGaussian. 3D Gaussians are optimized from multi-view images and SfM points. LightGaussian calculates each Gaussian’s global significance on training data, pruning those with the least significance. Next, SH coefficients are distilled into a compact format using synthesized pseudo-views. Finally, vector quantization, including codebook initialization and assignment, reduces model bandwidth.
@@
The LightGaussian framework is illustrated in Fig. 2. The 3D-GS model is trained on multi-view images, initialized from SfM point clouds, and expanded to millions of Gaussians to represent the scene comprehensively. Our pipeline then processes the 3D-GS model into a compact format using Gaussian Prune and Recovery to reduce the number of Gaussians, SH Distillation to eliminate redundant SHs while retaining key lighting information, and Vector Quantization to store Gaussians with lower bit-width.
@@ 3.1 Background: 3D Gaussian Splatting
where $\boldsymbol{\Sigma}$ can be decomposed into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$.
@@ 3.2 Gaussian Pruning & Recovery
Gaussian densification [16], which clones and refines the initial SfM point cloud, enhances small-scale geometry and detailed scene appearance by improving coverage. While this approach significantly boosts reconstruction quality, it increases the number of Gaussians from thousands to millions after optimization, resulting in substantial storage demands.Inspired by neural network pruning techniques [63] that remove less impactful neurons while preserving overall performance, we propose a tailored pruning strategy for 3D Gaussian Splatting to reduce over-parameterized points while maintaining accuracy. Identifying redundant yet recoverable Gaussians is essential to our approach. However, pruning Gaussians based on simple criteria (e.g., point opacity) risks degrading modeling performance, as it may eliminate essential scene details, as shown in Fig. 3.
@@
Inspired by Eq. 2, we go beyond using Gaussian opacity alone to assess significance by evaluating 3D Gaussians within the view frustum, projecting them onto the camera viewpoint for rendering. The initial significance score of each Gaussian (denoted as $\text{G}(\boldsymbol{X}_{j})$) can then be quantified based on its contribution to each pixel (ray, $r_{i}$) across all training views using the criteria $\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})$ whether they intersect or not. Consequently, we iterate over all training pixels to calculate the hit count of each Gaussian. The score is further refined by the adjusted 3D Gaussian’s volume $\large{\gamma}$$(\boldsymbol{\Sigma}_{j})$ and 3D Gaussian’s opacity $\sigma_{j}$, as they all contribute for the rendering formula. The volume calculation equation of the $j$-th 3D Guassian is V${(\boldsymbol{\Sigma}_{j})}=\frac{4}{3}\pi abc$, where $abc$ are the 3 dimensions of Scale ($\mathbf{S}$). We also consider the transmittance $\mathbf{T}$ is calculated with one subtract all the 3D Gaussian’s opacity the ray hit before the $j$-th 3D Gaussian, $\mathbf{T}={\prod_{i=1}^{j-1}}$ (1- $\sigma_{i}$). Finally, the global significance score can be summarized as:
$\text{GS}_{\text{j}}$ $=\sum_{i=1}^{MHW}\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})\cdot\sigma_{j}\cdot\mathbf{T}\cdot{\large{\gamma}}(\boldsymbol{\Sigma}_{j}),$ (3)
where $j$ is the Gaussian index, $i$ denotes a pixel, and $M$, $H$, and $W$ represent the number of training views, image height, and width, respectively. $\mathbbm{1}$ is an indicator function that determines whether a Gaussian intersects a given ray. However, using Gaussian volume alone tends to overemphasize background Gaussians, leading to excessive pruning of Gaussians modeling fine geometry. Thus, we propose a more adaptive approach to measuring volume dimensions:
@@
$\text{V}_{\text{norm}}$ $=\text{min}\left(\frac{\text{V}(\boldsymbol{\Sigma})}{\text{V}_{\text{max90}}},1\right).$
@@
We rank all Gaussians by their global significance scores to quantitatively guide pruning of the lower-ranked Gaussians. The remaining Gaussians are then jointly adapted by fine-tuning their attributes—without additional densification—to offset the minor loss from pruning. This adaptation is performed using photometric loss on the original training views, aligned with 3D-GS training for 5,000 iterations.
@@ 3.3 Distilling into Compact SHs
In the uncompressed Gaussian Splat representation, a substantial 81.3 percent of the feature dimension is occupied by Spherical Harmonics (SH) coefficients, requiring (45+3) floating-point values per splat. Directly reducing the SH degree can save disk space but also diminishes surface “shininess” and affects specular reflections.
To balance model size with scene quality, we introduce a knowledge distillation approach. Knowledge is distilled from a teacher model with full-degree SHs to a student model with truncated, lower-degree SHs. Supervision is based on the difference in predicted pixel intensities between the two models, with images synthesized at camera positions by sampling around each training view according to a Gaussian distribution:
$\mathcal{L}_{\text{distill}}$ $=\frac{1}{HW}\sum_{i=1}^{HW}\left\lVert\boldsymbol{C}_{\text{teacher}}(r_{i};[\mathbf{R}|\mathbf{t}])-\boldsymbol{C}_{\text{student}}(r_{i};[\mathbf{R}|\mathbf{t}])\right\rVert_{2}^{2}.$ (5)
@@
where $\mathbf{R}$ and $\mathbf{t}$ denote rendering camera rotation and position, $\mathbf{t}_{\text{pseudo}}$ and $\mathbf{t}_{\text{train}}$ represent the newly synthesized and training camera positions, respectively. $\mathcal{N}$ denotes a Gaussian distribution with mean 0 and variance $\sigma^{2}$, which is added to the original position to generate the new position.